# Uni-Instruct: One-step Diffusion Model through Unified Diffusion Divergence Instruction

### v2

@@ 3.3 How Uni-Instruct can Unify Previous Methods
[…]
@@
Complete proof is in Appendix A.4, which leverages integration by parts and Theorem 3.1.
@@ 3.4 Text-to-3D Generation using Uni-Instruct
Limitations of Uni-Instruct.
One of the major limitations of Uni-Instruct is that it needs an additional discriminator for density ratio estimation, which may bring more computational costs. Moreover, due to the complexity of the gradient formula, Uni-Instruct may result in bad performance with an improper choice of $f$, as its complex gradient formula is not as straightforward as some simpler existing methods like Diff-Instruct. We provide a detailed analysis in Appendix F.
4 Related Works
Diffusion Distillation
Diffusion distillation (luo2023comprehensive, ) focuses on reducing generation costs by transferring knowledge from teacher diffusion models to more efficient student models. It primarily includes three categories of methods: (1) Trajectory Distillation: These methods train student models to approximate the generation trajectory of diffusion models using fewer denoising steps. Approaches such as direct distillation (luhman2021knowledge, ; geng2023one, ) and progressive distillation (salimans2022progressive, ; meng2023distillation, ) aim to predict cleaner data from noisy inputs. Consistency-based methods (song2023consistency, ; kim2023consistency, ; song2023improved, ; liu2025scott, ; gu2023boot, ) instead minimize a self-consistency loss across intermediate steps. Most of these methods require access to real data samples for effective training. (2) Divergence Minimization (Distribution Matching): This line of work aims to align the distribution of the student model with that of the teacher. Adversarial training-based methods (xiao2021tackling, ; xu2024ufogen, ) typically require real data to perform distribution matching. Alternatively, several approaches minimize divergences like the KL divergence (e.g., Diff-Instruct (luo2023diff, ; yin2024one, )) or Fisher divergence (e.g., Score Identity Distillation (zhou2024score, ), Score Implicit Matching (luo2025one, )), and often do so without requiring real samples. Numerous improvements have been made to these two lines of work: DMD2 (yin2024improved, ) and SiDA (zhou2024adversarial, ) add real images during training, rapidly surpassing the teacher’s performance. $f$-distill (xu2025one, ) generalize KL divergence of DIff-Instruct into $f$-divergence and compared the affection of different divergences. Additionally, significant progress has been made toward scaling diffusion distillation for ultra-fast or even one-step text-to-image generation (luo2023latent, ; hoang2023swiftbrush, ; song2024sdxs, ; yin2024one, ; zhou2024long, ; yin2024improved, ). (3) Other Methods: Several alternative techniques that train the model from scratch have been proposed, including ReFlow (liu2022flow, ), Flow Matching Models (FMM) (boffi2024flow, ), which propose an ODE to model the diffusion process. Indcutive Moment Matching (zhou2025inductive, ) models the self-consistency of stochastic interpolants at different time steps. Consistency models (song2023consistency, ; song2023improved, ; kim2023consistency, ; geng2024consistency, ) impose consistency constraints on network outputs along the trajectory.
@@ 5.1 Benchmark Datasets Generation
Flow Rectified Flow (liu2022flow, ) 127 2.58
Flow Matching (lipman2022flow, ) 142 6.35
@@
Table 2: Class conditional ImageNet 64$\times$64 generation results. “Direct generation” and “Distillation” methods require one NFE, while the teacher uses 35 NFE. F.S. means from scratch. L.T. means resume and Longer Training.
@@
GAN BigGAN-deep (brock2018large, ) 1 4.06
StyleGAN-XL (sauer2022stylegan, ) 1 1.52
@@ 5.2 Ablation Studies
Figure 3: Prompt: A refined vase with artistic patterns. Left: ProlificDreamer; Right: UI+forward KL. Our vase demonstrate more diverse shapes as well as realistic patterns.
@@ 5.3 Text-to-3D Generation Using 2D Diffusion
In this subsection, we apply Uni-Instruct on text-to-3D generation. We re-implement the code base of ProlificDreamer (wang2023prolificdreamer, ) by adding an extra discriminator head to the output of the stable diffusion Unet’s encoder. We use FKL to distill the model for 400 epochs. Fig. 3 demonstrates the visual results from our 3D experiments. Uni-Instruct archives surprisingly decent 3D generation performances, with improved diversity and fidelity. Due to page limitations, we put detailed experiment settings and quantitative metrics in the Appendix E.

### v4

@@ 3.3 How Uni-Instruct can Unify Previous Methods
Suppose $W(t)=\int w(t)dt+C,W(0)=0$, the expression of Uni-Instruct with an extra weighting $W(t)$ is equivalent to $f$-distill:
@@
Complete proof is in Appendix B.4, which leverages integration by parts and Theorem 3.1.
@@ 4 Experiments
Table 3: Class conditional ImageNet $\times64$ generation results. “Direct generation” and “Distillation” methods require one NFE, while the teacher uses 35 NFE. F.S. means from scratch. L.T. means resume and Longer Training.
@@ 4.1 Benchmark Datasets Generation
4.2 Ablation Studies
We perform an ablation study on the techniques applied in our experiments. Table 5 ablates different components of our proposed method on CIFAR10, where we use an unconditional generator for all settings. For different divergences, we select three types: JKL, FKL, and RKL are divergences that only contains Grad(SiD), $\chi^{2}$ divergence’s gradient is only contributed by Grad(DI), Jensen-Shannon (JS) divergence has a gradient that contains both: $h_{\text{DI}}(\mathbf{x})\text{Grad(DI)}+h_{\text{SiD}}(\mathbf{x})\text{Grad(SiD)}$. Our result shows that JKL achieves the lowest FID value. Due to the numerical instability of the weightings, JS yields unsuccessful distillation results. As for the effect of GAN loss, we find that removing it still yields a decent result. Our integrated approach also surpasses the performance of using Uni-Instruct loss alone(without adding GAN loss), highlighting the effectiveness of combining expanded $f$-divergence with GAN losses. We also find that using a model trained with RKL Uni-Instruct (which recovers the SiD(zhou2024score, ) loss) as the initialization leads to better performances for all divergences.
Additional metrics evaluation and convergence analysis.
As is shown in the right table of Tab. 6, across different training iterations, our method consistently achieves lower FID scores than SiDA. This clearly indicates that our approach converges faster, reaching better generative quality with fewer iterations. Moreover, Tab. 6 and Tab. 7 compares the performance of Uni-Instruct and SiDA (zhou2024adversarial, ) across both CIFAR10 and ImageNet $64\times 64$ benchmarks on sFID, $\text{FD}_{\text{DINO}}$, inception score (IS), percision and recall. Uni-Instruct achieves strictly better results in 4 of 5 metrics on ImageNet $64\times 64$, the more complex and practically relevant benchmark. These gains are significant, particularly in $\text{FD}_{\text{dino}}$ and IS, which measure semantic alignment and perceptual quality/diversity, respectively.
The $18.7\%$ reduction in $\text{FD}_{\text{dino}}$ on ImageNet $64\times 64$ confirms Uni-Instruct’s advanced capability to preserve high-level semantic structures (e.g., object boundaries, textures, contextual relationships). This is critical for applications requiring fine-grained realism (e.g., medical imaging, autonomous driving). More importantly, Uni-Instruct’s gains widen significantly on ImageNet $64\times 64$ ($8.1\%$ IS, $18.7\%$ $\text{FD}_{\text{dino}}$) versus CIFAR10, proving its robustness for high resolution, semantically rich image synthesis. SiDA fails to maintain competitiveness under greater complexity.
Table 6: Further comparison between SiDA(zhou2024adversarial, ) and Uni-Instruct (forward KL). The left table is the performance evaluation on CIFAR10 unconditional generation. The right table compares the FID score along with the iterated k-images during training.
[…]
SiDA 1.71 132.72 10.32 0.670 0.624
Ours 1.66 129.30 10.30 0.671 0.626
$10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$
139.86 68.73 41.45 6.98 1.44
132.51 54.60 38.58 5.23 1.41
[…]
SiDA 1.68 111.26 10.28 0.678 0.632
Ours 1.68 108.89 10.29 0.679 0.629
[…]
1.98 74.86 59.28 0.562 0.653
2.01 60.86 64.11 0.561 0.658
Table 7: Further comparison between SiDA(zhou2024adversarial, ) and Uni-Instruct (forward KL). The left table is the performance evaluation on CIFAR10 conditional generation, while the right table is the performance evaluation on ImageNet $\times64$ generation.
@@ 4.2 Ablation Studies
Figure 3: Prompt: A refined vase with artistic patterns. Left: ProlificDreamer; Right: Uni-Instruct (forward KL). Our vase demonstrates more diverse shapes as well as realistic patterns.
@@ 4.3 Text-to-3D Generation Using 2D Diffusion
In this subsection, we apply Uni-Instruct on text-to-3D generation. We re-implement the code base of ProlificDreamer (wang2023prolificdreamer, ) by adding an extra discriminator head to the output of the stable diffusion Unet’s encoder. We use FKL to distill the model for 400 epochs. Fig. 3 demonstrates the visual results from our 3D experiments. Uni-Instruct archives surprisingly decent 3D generation performances, with improved diversity and fidelity. Due to page limitations, we put detailed experiment settings, quantitative metrics, and training algorithm in the Appendix E.